OpenStudy (anonymous):

Given the first term and the common difference of an arithmetic sequence find the first three terms and the explicit formula.

OpenStudy (anonymous):

\[a_{1}=28,d=10\]

OpenStudy (jhannybean):

First off, do you know the equation for an arithmetic sequence?/

OpenStudy (anonymous):

This? \[a_{n} = a + (n - 1)d.\]

OpenStudy (jhannybean):

Good :) Yes that is it.

OpenStudy (jhannybean):

Now we are looking for the first 3 terms, that can be written as \(n=3\)

OpenStudy (anonymous):

So I know that n=3 and d=10 Does a=28 in the equation?

OpenStudy (jhannybean):

Yes.

OpenStudy (jhannybean):

\(a=a_1=28\)

OpenStudy (anonymous):

So I have \[a_{3}=28+(3-1)10\] \[a_{3}=28+(2)10\] \[a_{3}=28+20\] \[a_{3}=48\]

OpenStudy (jhannybean):

Good :)

OpenStudy (jhannybean):

Now to find the explicit formula, we're basically writing the formula for \(a_3\) with all our terms without really solving for anything.

OpenStudy (anonymous):

Alright thanks! I get it now! It's been a while!

OpenStudy (jhannybean):

Haha yeah, for me as well!!